OpenStudy (anonymous):

How are the sin 30° and the cos 60° related? Explain using complete sentences.

OpenStudy (anonymous):

@eliassaab @jigglypuff314

OpenStudy (amoodarya):

if a+b=pi/2 =90 then sina = cos b cos a=sin b tan a= cot b

OpenStudy (anonymous):

If you have a 30-60-90 triangle, where one side is 1, the hypotenuse is 2, and the other side is sqrt(3) in length, then the values of sin30 and cos60 are equal. The 30 degree angle in this triangle is opposite the smallest side, of length 1, so taking sin30 gives you 1 over the hypotenuse (1/2). On the other hand, the side opposite to the 60 degree angle is of length sqrt3. Adjacent to the 60 degree angle (because cosine is opposite/hypotenuse) is the side of length 1, so taking cos60 gives you 1/2 (which is the same as sin30). Somebody better answer my question.

OpenStudy (anonymous):

sin(90-30) =cos 30 sin 60= cos30